The schedule below shows the quantities of shoes demanded at each income level in a community from 2020 to 2021:

 

Year

Income (R)

2020

5 000

1 000

Quantity (units) 2021

6 500

900


  1. Calculate the income elasticity of demand, using the arc method and interpret your result
  2. List any four factors that could result in the shoes having inelastic price elasticity of demand
  3. Explain the importance of understanding the relationship between price elasticity of demand and total revenue.
1
Expert's answer
2022-02-10T13:55:00-0500

Income Elasticity of Demand"=\frac{(\frac{∆InQuantityDemanded}{AverageQuantityDemanded)}}{\frac{∆InIncome}{Average Income}}"

"=\frac{\frac{6500-900}{3700}}{\frac{5000-1000}{3000}}"

"=0.75"

Therefore shoes are normal goods/Necessities since they have a positive elasticity of demand in relation to income.

2)

  • How the market is defined.
  • Availability of substitutes.
  • The time frame.
  • Share of income.

3) Firms need to understand price elasticity of demand as it greatly affects total revenue. Price and total revenue have a negative correlation when demand is elastic hence an increase in price will a decrease in total revenue.